OpenStudy (anonymous):

The graph of f is a subset of the domain of f. Is this always true or sometimes false?

OpenStudy (helder_edwin):

if \(f:A\to B\) is a function, then the domain of \(f\) is the whole of A.

OpenStudy (helder_edwin):

the graph of \(f:A\to B\) if a subset of \(A\times B\)

OpenStudy (helder_edwin):

so. there's no way u can compare the domain of a function with its graph.

OpenStudy (helder_edwin):

sorry *is a subset of ...

OpenStudy (anonymous):

But the question asks, whether it is always true or sometimes false.

OpenStudy (anonymous):

When is this ever true?

OpenStudy (helder_edwin):

never

OpenStudy (anonymous):

I have to choose either always true or sometimes false...

OpenStudy (anonymous):

did you leave?

OpenStudy (helder_edwin):

what subject is this for??

OpenStudy (anonymous):

multivariable calculus

OpenStudy (helder_edwin):

u could say "sometimes false"

OpenStudy (ash2326):

The graph of f is the subset of codomain so it's false so it's sometimes false

OpenStudy (anonymous):

Ok then could you explain to me when it is true?

OpenStudy (ash2326):

take example of simplest function f(x)=x domain= real numbers eg 1,2 -1 and so on range = real numbers so it's clearly subset of domain, isn't it?

OpenStudy (anonymous):

Ah... ok then what is an example when the statement is false? Sorry for so many questions. :|

OpenStudy (ash2326):

no problem. I'm here to help \[f(x)=-\sqrt {x}\] domain= positive real numbers range=negative real numbers clearly not a subset

OpenStudy (anonymous):

Ok so one last related question. "The domain of f is a subset of the range of f. Always true or sometimes false?" I am having great difficulty in seeing how domain can be related to the range in term of subsets.

OpenStudy (ash2326):

See my last post, is domain subset of range?

OpenStudy (anonymous):

Would this also be sometimes false then? Based on the previous examples you showed me?

OpenStudy (ash2326):

absolutely

OpenStudy (anonymous):

Thank you very very much. You have cleared htis up for me really fast. Also much thanks to helder_edwin too. :D This is an awesome website

OpenStudy (ash2326):

@hby0214 You're welcome